OpenStudy (anonymous):

Use the commutative law of multiplication to write an equivalent expression. (6 + x)y

OpenStudy (anonymous):

Commutative Law of Multiplication says: (a + b)y = y(a + b) I think so. Here a = 6 and b = x and now solve this...

OpenStudy (anonymous):

(6+x)y=y(6+x)

OpenStudy (anonymous):

Yes dear you have done it correctly..

OpenStudy (anonymous):

thank you